OpenStudy (anonymous):

Find y' by implicit differentiation: (9/x)+(9/y)=7. When I did this, I got -9y^2/x^2, but it is wrong. Not sure what I'm doing wrong

OpenStudy (anonymous):

did you solve \[-\frac{9}{x^2}-\frac{9}{y^2}y'=0\] for \(y'\) ?

OpenStudy (anonymous):

the 9's cancel

OpenStudy (anonymous):

so then does it become -y^2/x^2? I didn't cancel the 9's before.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh sweet thank you for helping!

OpenStudy (anonymous):

yw